# On irreducibility of six-dimensional compatible systems of $\mathbb{Q}$

### v1

@@ Abstract.
We study irreducibility of 6-dimensional strictly compatible systems of $\mathbb{Q}$ with distinct Hodge-Tate weights. We prove if one of the representations is Lie-irreducible, then all but finitely many of them are (Lie-)irreducible. If one of them is irreducible but not Lie-irreducible, we prove the same result under certain self-duality condition.
@@ 1. Introduction
Given a rational prime number $\ell$ and an elliptic curve $E$ over a number field $K$. The $\ell$-adic Tate module $\rho_{\ell}:=T_{\ell}(E)$ and $\ell$-torsion points $\overline{\rho}_{\ell}:=E[\ell]$ form classical examples of (2-dimensional) compatible systems and mod $\ell$ compatible systems. The following result asserts that they have uniform type of irreducible decompositions. Here (i) can be regarded as a weak version of Serre’s famous big image result in [Se72].
@@
    If $E$ has no complex multiplication over $\overline{K}$, then each $\rho_{\ell}$ is absolutely Lie-irreducible and $\overline{\rho_{\ell}}$ is absolutely irreducible for sufficiently large $\ell$.
@@
    If $E$ has potential complex multiplication but not over $K$, then $\{\rho_{\ell}\}$ is absolutely irreducible and induced by a 1-dimensional compatible system over a quadratic extension of $K$ after necessarily enlarging the coefficients. Also $\overline{\rho_{\ell}}$ is absolutely irreducible for sufficiently large $\ell$.
It is generally believed the above patterns are true for arbitrary semisimple compatible systems. More precisely we have the following 1.2. Note that in the above elliptic curve cases, the compatible systems are regular.
@@
    $\mathcal{M}$ can be decomposed as the direct sum of irreducible strictly compatible systems.
@@
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational regular strictly compatible system of $\mathbb{Q}$. Suppose some $\rho_{\lambda_{0}}$ is irreducible.
  1. (i)
    If $\rho_{\lambda_{0}}$ is Lie-irreducible, then $\rho_{\lambda}$ is (Lie-)irreducible for all but finitely many $\lambda$.
  2. (ii)
    If $\rho_{\lambda_{0}}$ is not Lie-irreducible but satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ with some character $\chi$, then $\rho_{\lambda}$ is irreducible for all but finitely many $\lambda$.
  3. (iii)
    If $\rho_{\lambda_{0}}$ satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ with some odd character $\chi$, then the semisimple reduction $\overline{\rho_{\lambda}}^{\text{ss}}$ is irreducible for all but finitely many $\lambda$.
@@
Let $\{\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational pure essentially self-dual and totally odd regular strictly compatible system of a totally real field $K$. Then there exists $r\in\mathbb{N}$ such that:
@@
We arrange the article as following. In section 2 we give necessary preliminaries for the proofs, including certain $\ell$-independence properties, big image results, potential automorphy theorem which is the main tool of the proof, and certain results on the semisimple reductions. In section 3 we prove the main results.
@@ 2.1. Compatible systems
  • $S$ is a finite set of primes in $K$ called exceptional set.
@@
    • $\rho_{\lambda}$ is unramified outside $S\cup S_{\lambda}$, where $S_{\lambda}$ is the primes of $K$ dividing the same rational prime as $\lambda$.
    • If $v\in S_{\lambda}$ then $\rho_{\lambda}|_{\mathrm{Gal}_{K_{v}}}$ is de Rham, and moreover is crystalline if $v\not\in S$.
@@
Under regularity condition, one can descend the coefficients of a strictly compatible system to $E_{\lambda}$ after possibly enlarging $E$:
@@
A strictly compatible system $\{\rho_{\lambda}\}$ is called pure of weight $\omega$, if for each $v\not\in S$, for each root $\alpha$ of $P_{v}(T)$ in $\overline{E}$ and for each embedding $i:\overline{E}\hookrightarrow\mathbb{C}$, one has $|i\alpha|^{2}=(\#\kappa(v))^{\omega}$, where $\kappa(v)$ is the residue field at $v$.
@@
  • $S$ is a finite set of primes in $K$ called exceptional set.
@@
    $\rho_{\lambda}$ is unramified outside $S\cup S_{\lambda}$, where $S_{\lambda}$ is the primes of $K$ dividing the same rational prime as $\lambda$.
@@ 2.2. $\lambda$-independence
Given a $\lambda$-adic semisimple continuous Galois representation
$\rho_{\lambda}:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{E_{\lambda}})$
Its algebraic monodromy group $\textbf{G}_{\lambda}$ is defined to be the Zariski closure of the image inside algebraic group $\text{GL}_{n,\overline{E_{\lambda}}}$.
@@
Given a $\lambda$-adic Galois representation $\rho$. Assume $\sigma$ is a subquotient of $\rho$. Then the algebraic monodromy groups of $\rho$ and $\rho\oplus\sigma$ coincide.
@@
The algebraic monodromy group of $\rho$ is the fundamental group of the Tannakian category generated by $\rho$ inside the category of all $\lambda$-adic Galois representations. The result follows from the fact $\rho$ and $\rho\oplus\sigma$ generate the same Tannakian category. ∎
Due to semisimplicity, the identity component $\textbf{G}_{\lambda}^{\circ}$ is a reductive group. We write $\textbf{G}_{\lambda}^{\text{der}}$ to be the derived subgroup of $\textbf{G}_{\lambda}^{\circ}$, which is semisimple.
@@
Let $F$ be a field and $\textbf{G}\subseteq\mathrm{GL}_{n,F}$ a reductive subgroup.
@@
    Dnote by $\textbf{T}^{\prime}$ the maximal torus of $\textbf{G}^{\mathrm{der}}\times\overline{F}$. Then the formal bi-character of G is the conjugacy class of the chain $^{\prime}\subseteq\textbf{T}$ T in $\mathrm{GL}_{n,\overline{F}}$.
@@
    Let $\{F_{i}\}$ be a family of fields and $\{\textbf{G}_{i}\subseteq\mathrm{GL}_{n,F_{i}}\}$ be a family of reductive groups. We say they have same formal bi-character if there exists a chain of subtorus $\textbf{T}_{\mathbb{Z}}^{\prime}\subseteq\textbf{T}_{\mathbb{Z}}\subseteq\mathrm{GL}_{n,\mathbb{Z}}$ such that the chain $\textbf{T}_{\mathbb{Z}}^{\prime}\times\overline{F}_{i}\subseteq\textbf{T}_{\mathbb{Z}}\times\overline{F}_{i}$ is contained in formal bi-character of $\textbf{G}_{i}$ for each $i$.
We have the following standard $\lambda$-independence results on algebraic monodromy groups.
@@
Denote by $(\overline{\rho}^{\text{ss}},\overline{V}^{\text{ss}})$ the semisimple reduction of a $\lambda$-adic Galois representation $(\rho,V)$. For the purpose of investigating residual irreducibility, we need the following concept.
@@
We mention a recent result on $\ell$-independence of component groups of algebraic envelope groups which is the analogues of Serre’s result 2.9(i). Given a smooth projective variety $X$ over some number field $K$, then the semisimplification of $\ell$-adic cohomology
@@ 2.3. Lie-irreduciblility
A $\lambda$-adic continuous Galois representation $\rho:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{E_{\lambda}})$ is called Lie-irreducible if for any finite extension $L/K$, the restriction $\rho|_{\mathrm{Gal}_{L}}$ is irreducible.
@@
where each $\sigma_{i}$ is Lie-irreducible of $K_{i}$. Denote by $\text{Spl}(\rho)$ the set of primes in $K$ which have a split factor in at least one $K_{i}$. Then the usual formula for the trace of an induced representation shows $\text{Spl}(\rho)$ is contained in the set of trace zero (under $\rho$) primes of $K$. [Pa19, Proposition 3.4.9.(1)] shows the Dirichlet density of the two sets coincide. Hence by compatibility of the system we obtain:
@@
2.4. Essential self-duality
@@
Our condition $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi$ in 1.3.(ii) is slightly general than essential self-duality. If a representation satisfies this condition, again we call $\chi$ the similitude character. Note that the similitude character may not be unique.
@@
A Galois character $\chi$ of a totally real field is called totally odd (respectively, totally even), if for any complex conjugation $c$ one has $\chi(c)=-1$ (respectively, $\chi(c)=1$).
@@
Given an $E$-rational Serre compatible system $\{\rho_{\lambda}\}$ and one extends the scalars of each $\rho_{\lambda}$ to $\overline{E}_{\lambda}$. If some $\rho_{\lambda_{0}}$ satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi_{\lambda_{0}}$. Then each $\rho_{\lambda}$ satisfies $\rho_{\lambda}\cong\rho_{\lambda}^{\vee}\otimes\chi_{\lambda}$ for some character $\chi_{\lambda}$. Moreover if the simititude character $\chi_{\lambda_{0}}$ is totally odd (resp. totally even), then one can choose each simititude character $\chi_{\lambda}$ to be totally odd (resp. totally even).
@@ 2.5. Automorphic Galois representations
The following result shows for 2-dimensional representations of totally real fields, one can rule out even cases when $\lambda$ is sufficiently large.
@@
Then the pair $(\rho,\det\rho)$ is essentially self-dual and odd.
In particular, above results shows certain low dimensional subrepresentations of strictly compatible systems fit into strictly compatible systems.
@@
consists of only 1 or 2 dimensional regular irreducible conponents. Then the compatible system $\{\rho_{\lambda}\}$ can be written as direct sum of some 1 or 2 dimensional irreducible strictly compatible systems.
@@ 2.6. Semisimple reduction
We use the following result in the proof. See also [PSW18, Corollary 1.3] to remove regularity condition.
@@ 3.1. The non Lie-irreducible case
$\rho_{\lambda_{0}}=\text{Ind}^{\mathbb{Q}}_{K}\sigma$
@@ 3.1.1. $[K:\mathbb{Q}]=6$
In such case $\sigma$ is an algebraic $\lambda$-adic character. By class field theory after enlarging $E$, this $\sigma$ extends to an $E$-rational compatible system $\{\sigma_{\lambda}\}$. Then $\{\text{Ind}^{\mathbb{Q}}_{K}\sigma_{\lambda}\}$ also is an $E$-rational compatible system. By semisimplicity each $\rho_{\lambda}$ is isomorphic to $\text{Ind}^{\mathbb{Q}}_{K}\sigma_{\lambda}$. Then the irreducibility of all $\rho_{\lambda}$ comes from Mackey’s irreducibility criterion and regularity. In particular this $\{\rho_{\lambda}\}$ is induced by a one-dimensional strictly compatible system of $F$.
@@ 3.1.2. $[K:\mathbb{Q}]=3$
2.16 rules out the situation that some irreducible component of $\rho_{\lambda}$ is Lie-irreducible. We also assume we are not in the case that each irreducible component of $\rho_{\lambda}$ is induced by some character, since otherwise by regularity and class field theory, we can use the same method as in 3.1.1 to show $\rho_{\lambda_{0}}$ is reducible. The only remaining case is:
@@
We have shown each $\rho_{\lambda}$ is induced by some 2-dimensional $\sigma_{\lambda}$ of some cubic field $K_{\lambda}$. If $K/\mathbb{Q}$ is Galois, then all $K_{\lambda}$ coincide. This is because when $K/\mathbb{Q}$ is Galois, restricting $\rho_{\lambda_{0}}$ to $K$ has dimensional type $2+2+2$ since the Galois group acts transitively on the set of irreducible subrepresentations. But if some $K_{\lambda}\not=K$, restricting $\mathrm{Ind}_{K_{\lambda}}^{\mathbb{Q}}\sigma_{\lambda}$ to $K$ still is irreducible and induced by a 2-dimensional Lie-irreducible representation of $KK_{\lambda}$. This contradicts with 2.15. Moreover, as $K$ is also totally real, restrict $\{\rho_{\lambda}\}$ to $K$ then apply 2.23.(i), one obtains some $\sigma_{\lambda}$ fits into a compatible system. Hence by semisimplicity, one writes $\{\rho_{\lambda}\}$ as the induced representations of a 2-dimensional Lie-irreducible strictly compatible system of $K$.

### v3

@@ Abstract.
We study irreducibility of 6-dimensional strictly compatible systems of $\mathbb{Q}$ with distinct Hodge-Tate weights. We prove if one of the representations is irreducible, then all but finitely many of them are irreducible.
@@ 1. Introduction
Given an elliptic curve $E$ over a number field $K$. When $\ell$ varies over rational primes, the rational $\ell$-adic Tate modules $\rho_{\ell}:=T_{\ell}(E)\otimes\mathbb{Q}_{\ell}$ and $\ell$-torsion points $\overline{\rho}_{\ell}:=E[\ell]$ form classical examples of (2-dimensional) compatible system and mod $\ell$ compatible system. The following result asserts that they have uniform type of irreducible decompositions. Here (i) can be regarded as a weak version of Serre’s famous big image result in [Se72].
@@
    If $E$ has no complex multiplication over $\overline{K}$, then each $\rho_{\ell}$ is absolutely irreducible and Lie-irreducible for all $\ell$ and $\overline{\rho}_{\ell}$ is absolutely irreducible for sufficiently large $\ell$.
@@
    If $E$ has potential complex multiplication but not over $K$, then $\{\rho_{\ell}\}$ is absolutely irreducible and induced by a 1-dimensional compatible system of a quadratic extension of $K$ after necessarily enlarging the coefficients. Also $\overline{\rho}_{\ell}$ is absolutely irreducible for sufficiently large $\ell$.
It is generally believed the above patterns are true for arbitrary semisimple compatible systems. More precisely we have the following 1.2. We call a compatible system irreducible (resp. Lie-irreducible) if each representation is irreducible (resp. Lie-irreducible). Note also that in above elliptic curve cases, the compatible systems are regular.
@@
    $\mathcal{M}$ can be decomposed as a direct sum of irreducible strictly compatible systems.
@@
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational regular strictly compatible system of $\mathbb{Q}$. Suppose some $\rho_{\lambda_{0}}$ is irreducible, then $\rho_{\lambda}$ is irreducible for all but finitely many $\lambda$.
@@
Let $\{\rho_{\lambda}:\mathrm{Gal}_{\mathbb{Q}}\to\text{GL}_{6}(\overline{E_{\lambda}})\}$ be a 6-dimensional $E$-rational pure essentially self-dual and totally odd regular strictly compatible system of rational numbers $\mathbb{Q}$. Then there exists $r\in\mathbb{N}$ such that:
@@
We arrange the article as following. In section 2 we give necessary preliminaries for the proofs, including certain $\ell$-independence properties, big image results, potential automorphy theorem which is the main tool of the proof, and certain results on the semisimple reductions. In section 3 we prove the main results. To prove 1.3 we consider whether $\rho_{\lambda_{0}}$ is Lie-irreducible or not, which are treated in section 3.1 and 3.2 separately. Section 3.3 proves 1.4.
@@ 2.1. Compatible systems
  • $S$ is a finite set of primes of $K$ called exceptional set.
@@
    • $\rho_{\lambda}$ is unramified outside $S\cup S_{\lambda}$, where $S_{\lambda}$ is the primes of $K$ that divide the same rational prime as $\lambda$.
    • If $v\in S_{\lambda}$ then $\rho_{\lambda}|_{\mathrm{Gal}_{K_{v}}}$ is de Rham.
    Moreover, $\rho_{\lambda}|_{\mathrm{Gal}_{K_{v}}}$ is crystalline if $v\in S_{\lambda}$ and $v\not\in S$.
@@
Under regularity condition, one can descend the coefficients of a strictly compatible system to $E_{\lambda}$ after enlarging $E$:
@@
An $E$-rational strictly compatible system $\{\rho_{\lambda}\}$ is called pure of weight $\omega$, if for each $v\not\in S$, for each root $\alpha$ of $P_{v}(T)$ in $\overline{E}$ and for each embedding $i:\overline{E}\hookrightarrow\mathbb{C}$, one has $|i\alpha|^{2}=(\#\kappa(v))^{\omega}$, where $\kappa(v)$ is the residue field at $v$.
@@
  • $S$ is a finite set of primes of $K$ called exceptional set.
@@
    $\rho_{\lambda}$ is unramified outside $S\cup S_{\lambda}$, where $S_{\lambda}$ is the primes of $K$ that divide the same rational prime as $\lambda$.
@@ 2.2. $\lambda$-independence
Given a semisimple $\ell$-adic Galois representation
$\rho:\mathrm{Gal}_{K}\to\text{GL}_{n}(E_{\lambda})$
The algebraic monodromy group G of $\rho$ is defined to be the Zariski closure of the image inside algebraic group $\text{GL}_{n,E_{\lambda}}$.
@@
Given a semisimple $\ell$-adic Galois representation $\rho$. Assume $\sigma$ is a subquotient of $\rho$. Then the algebraic monodromy groups of $\rho$ and $\rho\oplus\sigma$ coincide.
@@
The algebraic monodromy group of $\rho$ is the fundamental group of the Tannakian category generated by $\rho$ inside the category of all $\ell$-adic Galois representations. The result follows from the fact that $\rho$ and $\rho\oplus\sigma$ generate the same Tannakian category. ∎
Due to semisimplicity, the identity component $\textbf{G}^{\circ}$ is a reductive group. We write $\textbf{G}^{\text{der}}$ to be the derived subgroup of $\textbf{G}^{\circ}$, which is semisimple. To describe $\lambda$-independent properties of compatible systems we need following concepts.
@@
Let $F$ be a field and $\textbf{G}\subseteq\mathrm{GL}_{n,F}$ be a reductive subgroup.
@@
    Denote by $\textbf{T}^{\prime}$ the maximal torus of $\textbf{G}^{\mathrm{der}}\times\overline{F}$. Then the formal bi-character of G is the conjugacy class of the chain $^{\prime}\subseteq\textbf{T}$ T in $\mathrm{GL}_{n,\overline{F}}$.
@@
    Given two fields $F_{1},F_{2}$ and two reductive groups $\textbf{G}_{i}\subseteq\mathrm{GL}_{n_{i},F_{i}},i=1,2$. We say they have same formal character (resp. formal bi-character), if $n_{1}=n_{2}=n$ and there exists a split $\mathbb{Z}$-subtorus $\textbf{T}_{\mathbb{Z}}\subseteq\mathrm{GL}_{n,\mathbb{Z}}$ (resp. a chain of split $\mathbb{Z}$-subtori $\textbf{T}_{\mathbb{Z}}^{\prime}\subseteq\textbf{T}_{\mathbb{Z}}\subseteq\mathrm{GL}_{n,\mathbb{Z}}$) such that $\textbf{T}_{\mathbb{Z}}\times\overline{F_{i}}$ (resp. $\textbf{T}_{\mathbb{Z}}^{\prime}\times\overline{F}_{i}\subseteq\textbf{T}_{\mathbb{Z}}\times\overline{F}_{i}$) is contained in formal character (resp. formal bi-character) of $\textbf{G}_{i}$ for each $i$. This defines an equivalence relation on formal characters (resp. formal bi-characters) of reductive groups over different fields.
  4. (iv)
    Let $\{F_{i}\}$ be a family of fields and $\{\textbf{G}_{i}\subseteq\mathrm{GL}_{n,F_{i}}\}$ be a family of reductive groups. We say they have same formal character (resp. same formal bi-character) if they belong to the same class under the equivalence relation in (iii). We say they have bounded formal characters (resp. bounded formal bi-characters) if they belong to finitely many classes under the equivalence relation in (iii).
We have following standard $\lambda$-independence results on algebraic monodromy groups.
@@
Denote by $(\overline{\rho}^{\text{ss}},\overline{V}^{\text{ss}})$ the semisimple reduction of a $\lambda$-adic Galois representation $(\rho,V)$. For the purpose of investigating residual irreducibility, we need following concept.
@@
We mention a recent result on $\ell$-independence of component groups of algebraic envelopes which is the analogues of Serre’s result 2.9(i). Given a smooth projective variety $X$ over some number field $K$, then the semisimplification of $\ell$-adic cohomology
@@ 2.3. Lie-irreduciblility
A continuous Galois representation $\rho:\mathrm{Gal}_{K}\to\text{GL}_{n}(\overline{E_{\lambda}})$ is called Lie-irreducible if for any finite extension $L/K$, the restriction $\rho|_{\mathrm{Gal}_{L}}$ is irreducible.
@@
where each $\sigma_{i}$ is Lie-irreducible of $K_{i}$. Denote by $\text{Spl}(\rho)$ the set of primes in $K$ which have a split factor in at least one $K_{i}$. Then the usual formula for the trace of an induced representation shows that the set of trace non-zero (under $\rho$) primes of $K$ is contained in $\text{Spl}(\rho)$. [Pa19, Proposition 3.4.9.(1)] shows the Dirichlet density of the two sets coincide. Hence by compatibility of the system we obtain:
@@
2.4. Essential self-duality and oddness
@@
Note that the similitude character may not be unique.
@@
A Galois character $\chi$ of a totally real field is called totally odd (resp. totally even), if for any complex conjugation $c$ one has $\chi(c)=-1$ (resp. $\chi(c)=1$).
@@
Given an $E$-rational Serre compatible system $\{\rho_{\lambda}\}$ and one extends the scalars of each $\rho_{\lambda}$ to $\overline{E}_{\lambda}$. If some $\rho_{\lambda_{0}}$ satisfies $\rho_{\lambda_{0}}\cong\rho_{\lambda_{0}}^{\vee}\otimes\chi_{\lambda_{0}}$. Then each $\rho_{\lambda}$ satisfies $\rho_{\lambda}\cong\rho_{\lambda}^{\vee}\otimes\chi_{\lambda}$ for some character $\chi_{\lambda}$. Moreover if the similitude character $\chi_{\lambda_{0}}$ is totally odd (resp. totally even), then one can choose each similitude character $\chi_{\lambda}$ to be totally odd (resp. totally even).
@@ 2.5. Automorphic Galois representations
The following result shows under certain minor conditions, for 2-dimensional representations of totally real fields, the representation is totally odd when $\lambda$ is sufficiently large.
@@
Then the pair $(\rho,\det\rho)$ is essentially self-dual and totally odd.
In particular, above results show certain low dimensional subrepresentations of strictly compatible systems fit into strictly compatible systems.
@@
consists of only 1 or 2 dimensional regular irreducible conponents. Then the compatible system $\{\rho_{\lambda}\}$ can be written as a direct sum of some 1 or 2 dimensional irreducible strictly compatible systems.
@@ 2.6. Semisimple reduction
We use the following result in the proof.
@@ 3.1. The non Lie-irreducible case
(1) $\rho_{\lambda_{0}}=\text{Ind}^{\mathbb{Q}}_{K}\sigma$
@@ 3.1.1. $[K:\mathbb{Q}]=6$
In such case $\sigma$ is an algebraic $\lambda$-adic character. By class field theory after enlarging $E$, this $\sigma$ extends to an $E$-rational compatible system $\{\sigma_{\lambda}\}$. Then $\{\text{Ind}^{\mathbb{Q}}_{K}\sigma_{\lambda}\}$ also is an $E$-rational compatible system. By semisimplicity each $\rho_{\lambda}$ is isomorphic to $\text{Ind}^{\mathbb{Q}}_{K}\sigma_{\lambda}$. Then the irreducibility of all $\rho_{\lambda}$ comes from Mackey’s irreducibility criterion and regularity. In particular this $\{\rho_{\lambda}\}$ is induced by a one-dimensional strictly compatible system of $K$.
@@ 3.1.2. $[K:\mathbb{Q}]=3$
2.16 rules out the situation that some irreducible component of $\rho_{\lambda}$ is Lie-irreducible. We also assume we are not in the case that each irreducible component of $\rho_{\lambda}$ is induced by some character, since otherwise by semisimplicity and class field theory, we can use the same method as in 3.1.1 to show $\rho_{\lambda_{0}}$ is reducible. The only remaining case is:
@@
We have shown each $\rho_{\lambda}$ is induced by some 2-dimensional $\sigma_{\lambda}$ of some cubic field $K_{\lambda}$. In fact we can show one can take $K_{\lambda}=K$ in each $\rho_{\lambda}$ up to isomorphsim. Moreover if $K/\mathbb{Q}$ is Galois, then the compatible system can be written as induced representations of a 2-dimensional Lie-irreducible compatible system of $K$.
Proof.
We first show one can take $K_{\lambda}=K$ in each $\rho_{\lambda}$ up to isomorphsim. Suppose first $K/\mathbb{Q}$ is Galois. The restriction of $\rho_{\lambda_{0}}$ to $K$ has dimensional type $2+2+2$ since the Galois group acts transitively on the set of irreducible subrepresentations. If some $K_{\lambda}\not=K$, the restriction of $\mathrm{Ind}_{K_{\lambda}}^{\mathbb{Q}}\sigma_{\lambda}$ to $K$ still is irreducible and is induced by a 2-dimensional Lie-irreducible representation of $KK_{\lambda}$. This contradicts with 2.16.
Suppose now $K/\mathbb{Q}$ is not Galois. Denote $G=\text{Gal}_{\mathbb{Q}}$, $H=\text{Gal}_{K}$ and $H_{\lambda}=\text{Gal}_{K_{\lambda}}$. Consider the compatible system $\{\text{Res}_{K}\rho_{\lambda}\}$, more precisely:
$\text{Res}_{K}\rho_{\lambda}=\text{Ind}^{K}_{KK_{\lambda}}\sigma_{\lambda}|_{KK_{\lambda}}\oplus\left(\bigoplus_{1\not=s\in H\backslash G/H_{\lambda}}\text{Ind}^{H}_{sH_{\lambda}s^{-1}\cap H}\sigma_{\lambda}^{s}\right)$
where $\sigma_{\lambda}^{s}(x)=\sigma_{\lambda}(s^{-1}xs)$ for $x\in sH_{\lambda}s^{-1}\cap H$. For $\lambda=\lambda_{0}$ this restriction has a Lie-irreducible component $\sigma_{\lambda}$, hence by 2.16, $\text{Res}_{K}\rho_{\lambda}$ cannot be induced type. This rules out $[KK_{\lambda}:K]=3$. Moreover assume $K\not=K_{\lambda}$ then $[KK_{\lambda}:K]\not=1$. Hence by dimensional reason there exists $s\in G$ such that $sH_{\lambda}s^{-1}=H$. This shows $K$ and $K_{\lambda}$ are isomorphic. Hence for those $K_{\lambda}\not=K$ we can modify $\rho_{\lambda}$ by:
$\text{Ind}^{G}_{H_{\lambda}}\sigma_{\lambda}\cong\text{Ind}^{G}_{sH_{\lambda}s^{-1}}\sigma_{\lambda}^{s}=\text{Ind}^{G}_{H}\sigma_{\lambda}^{s}$
where $\sigma_{\lambda}^{s}(x)=\sigma_{\lambda}(s^{-1}xs)$. Hence we can choose $K_{\lambda}=K$ in each $\rho_{\lambda}$ up to isomorphism.
Finally, when $K/\mathbb{Q}$ is cubic Galois, $K$ also is totally real. Restrict $\{\rho_{\lambda}\}$ to $K$ then apply 2.23.(i), one obtains some $\sigma_{\lambda}$ fits into a compatible system. Hence by semisimplicity, one writes $\{\rho_{\lambda}\}$ as induced representations of a 2-dimensional Lie-irreducible strictly compatible system of $K$. ∎